Doreen and Irena plan to leave their houses at the same time, roller
blade towards each other, and meet for lunch after 2 hours on the road.
Doreen can maintain a speed of 2 miles per hour, which is 40% of Irena's
speed. If they meet exactly as planned, what is the distance between
their houses?
I don't know how to solve using liner equations?

Answer by Theo(13342)

rate * time = distance is the basic formula to use for this type of problem.

they are both skating for 2 hours when they meet.

doreen skates at 2 miles per hour.

this is 40% of irena's rate.

if we let x = irena's speed, then 2 = .40 * x

solve for x to get x = 2 / .40 = 5 miles per hour.

doreen skates at 2 miles per hour.

irena skates at 5 miles per hour.

2/5 = .4 = 40% so we got the speeds right for doreen and irena.

rate * time = distance is the basic formula to use.

in 2 hours, doreen skates 2 * 2 = 4 miles towards irena.

in 2 hours, irena skates 5 * 2 = 10 miles towards doreen.

the distance between them has to be 14 miles.